What i need done is tough!!!!Its like word problems and we make a table and stuff!!!!! What i need done is just an example of like: x^3 + 2x^2 + x + 12
--------------------
x + 3
I need a table and please show the work of the equation!!!!

Answer by stanbon(48516)

Fredique is 2 years older than her brother. 12 years ago she was twice as old as he was.
--------------
Label two rows as "fredique" and "brother"
Draw two columns to the right of the names
Label the columns "Now" and "12 yrs. ago.
-----------
In the Now column put x as the age of brother and x+2 as the age of Fredique.
In the 12 yrs ago column place "x-12" in the brother row; put x+2-12= x-10
in the fredrique row.
Below the table write EQUATION:
Using the data in the "12 yrs from now" symbolize the statement "Fredrique
is twice as old as her brother"
x-10 = 2(x-12)
x-10 = 2x-24
x=14 (brother's age now)
x+2=16 (Fredrique age now)
---------
Cheers,
Stan H>